# Electron-phonon coupling in magnetic materials using the local spin density approximation

### v1

@@ III Methods
DFT calculations are performed using Quantum ESPRESSO (QE[60, 61] version 7.4 and Abinit version 10.2.5 [77, 56, 57, 59]. Scalar-relativistic norm-conserving pseudopotentials [78] from the standard accuracy table of the PseudoDojo project [79] are used. We use the local density approximation (LDA) [80] and the generalized gradient approximation (GGA) with the Pardew Burke Ernzerhof (PBE) functional [81]. The interpolated phonon dispersions and dynamical matrices are obtained with DFPT [42, 44, 43, 45]. The maximally localized Wannier functions are obtained using the wannier90 package version 3.1 [82]. Direct evaluation of spin-resolved EPME is obtained by modifying QE to print the values. We provide the modified code on the Materials Cloud Archive [83] along with all data to reproduce the results presented here. The interpolation of electronic band structures, phonon band structures, and electron phonon matrix elements, as well as the calculation of transport properties and electron-phonon coupling matrix elements are performed by modifying the EPW package version 5.9 [64]. This feature will be released in a future EPW version and we also provide the code on the Materials Cloud Archive [83].
The cutoff energy for the plane wave expansion, $\bf k$-meshes for the electrons, and the $\bf q$-meshes for the phonons are selected so that an arbitrary increase in those parameters produces a total energy change of less than 1 meV per unit cell. This corresponds to 100 Ry for both Fe and Ni, as well as a $\bf k$-point grid of 243 and a $\bf q$-point grid of 83 for both Ni and Fe. Structural relaxations are performed with a convergence criterion of $\Delta E<10^{-12}$ Ry, and with a pressure of less than 1 bar. In the case of phonons, the convergence criteria in order to obtain accurate force constants was tr2_ph $10^{-20}$ for QE while for Abinit the convergence criteria is on the energy $\Delta E<10^{-20}$ Ha. We converge the resistivity below 1% above 150 K using EPW, and validate the results with an in-house developed version of the ElectronPhonon.jl package [84, 85], which relies on the real-space EPME computed with the EPW code [63, 64] that we extend to the collinear magnetic case for the purpose of this work. Excellent agreement is obtained between the two codes, see Supplemental Material S1 [83] for more information.
@@ IV Electron-phonon coupling with magnetism
In this work, we study BCC Fe and FCC Ni, both in their ferromagnetic configuration using the collinear approximation. We also consider non-magnetic Pb and verify that switching collinear magnetism produced the same result as non-magnetic calculations, as described in Supplementary Material S2 [83].
@@ IV.1 Verification of the direct evaluation of the electron-phonon coupling
To verify our implementation, we compare the QE results with those obtained with Abinit. Although both codes match well in non-magnetic compounds [65, 66, 67], no validation study had yet been performed in the magnetic case. To this end, we compare basic quantities including total energy, total magnetization and phonon frequencies while keeping the computational parameters as close as possible. We find an excellent level of agreement between the two codes that is similar to the non-magnetic case, see Supplementary Material S3 [83].
@@
We compare the absolute value of the deformation potentials between the two codes. To achieve this, we note that QE performs an average over all degenerate states in electronic and phonon band indices. We therefore applied the same average to the Abinit results and present a comparison for ferromagnetic Fe in Fig. 1 between the two codes for $\mathbf{k}=\bm{\Gamma}$ and bands $m=n=5$. We find a good agreement between the two codes with a maximum error of less than 0.02 meV/bohr. We provide the same comparison for Ni in Supplemental Material S4 [83] where we find an error below 0.01 meV/bohr.
@@ IV.2 Verification of the electron phonon interpolation
Figure 3: Deformation potential for the second to sixth lowest electronic bands in Fig. 2(a) and 2(b) for (a) Fe and (b) Ni at $\mathbf{k}=\bm{\Gamma}$ for $\uparrow$ (red) and $\downarrow$ (blue). Solid dots are direct evaluation of the deformation potential with Quantum ESPRESSO. The lines are the EPW interpolated deformation potential from $4^{3}$, $6^{3}$, $8^{3}$, $10^{3}$ and $12^{3}$ $\mathbf{k/q}$ homogeneous coarse grids with a color gradient from lighter to darker representing coarser to denser grids. Both spin channels are close in Ni while a larger difference is observed for Fe.
We first compare in Fig. 2(a) and 2(b) the direct and Wannier-interpolated electronic band structures of ferromagnetic Fe and Ni respectively, obtaining an energy mismatch of less than 0.1 meV for all the displayed $\mathbf{k}$-points and bands. This shows that both spin channels are correctly reproduced by the interpolation procedure. We then compute in Fig. 2(c) and 2(d) the phonon dispersion of ferromagnetic Fe and Ni respectively using LDA and PBE. Overall, LDA overestimates by about 10% the phonon frequency with respect to neutron scattering experiments [86, 87]. Using the PBE functional yields a much better agreement, in line with a previous study [88]. Moreover, we validate the interpolated electron-phonon deformation potential by comparing to direct calculations in Fig. 3, for electronic bands crossing the Fermi level and at the electronic zone center. We show the two spin channels for Fe and Ni, and find that convergence is achieved for $\times12$$\times$12 $\mathbf{k}/\mathbf{q}$ coarse grids in both cases. For such grids, we find excellent agreement between interpolated and independently calculated reference direct calculations. In addition, we also compare the EPW interpolation with the Abinit interpolation finding similar results; see Supplementary Material S4 [83]. Finally, we verify that the Hamiltonian, dynamical matrices, and EPME all decay rapidly in real space; see Supplementary Material S5 [83].
@@ V Physical properties and electron-phonon quantities
Compound Magnetism $\lambda$ $\lambda^{\uparrow}$ $\lambda^{\downarrow}$ $\omega_{\mathrm{log}}[\rm K]$ $T_{\rm c}^{\rm AD}[\rm K]$
Fe (BCC) FM 0.2961 0.2180 0.0781 311 0.02
Ni (FCC) FM 0.2767 0.0040 0.2727 291 0.01
Table 1: Total and spin-resolved electron phonon coupling strength $\lambda$ for ferromagnetic (FM) Fe and Ni. Evaluated characteristic frequency $\omega_{\mathrm{log}}$ and Allen-Dynes critical temperature $T_{\mathrm{c}}^{\mathrm{AD}}$ with $\mu^{*}=0.13$.
@@
We also compute the Eliashberg spectral functions and show the result in Fig. 5. We see in Fig. 5(a) that the Fe spin majority channel is the dominant contribution for most of the energy range. In contrast, for Ni the spin minority dominates throughout and the spin majority is negligible, see Fig.  5(b). This important difference is due to the density of states at the fermi level in Ni, see Fig. 4. Using the $\alpha^{2}F^{\sigma}(\omega)$, we evaluate the $\omega_{\mathrm{log}}$ and compute the $T_{\rm c}$ with the Allen-Dynes formula. In doing so, we need to provide a reasonable value for the semi-empirical Coulomb potential $\mu^{*}$. Typical values range between 0.1 and 0.2 in the literature  [89, 90, 91] and we choose $\mu^{*}=0.13$ here. We find critical temperatures of 0.02 K and 0.01 K respectively for Fe and Ni (see Table 1), showing the absence of Eliashberg superconductivity even if the pairing would not be affected by the internal magnetization. Performing a scan from $\mu^{*}=0.1-0.2$ we find that in the case of Fe, the $T_{c}$ ranges from 0.13 to 0 K while in the case of Ni $T_{c}$ ranges from 0.06 to 0 K. We note that magnetism could lead to a larger effective $\mu^{*}$ since it needs to account for the Coulomb potential and spin fluctuations [89], which would further decrease $T_{\rm c}$, even within Eliashberg theory.
The results obtained on the $T_{\rm c}$ of Fe and Ni are compatible with ambient pressure experiments, which never show a superconducting transition. Fe may show a superconducting transition at very low temperature $T_{c}\leq 2$ K under high pressure [92, 93].
@@ V.1 Resistivity of magnetic metals
We now focus on the resistivity of metals and solve the BTE for both spin channels in ferromagnetic Fe and Ni. We note that when converging the resistivity and in general electron-phonon related quantities, it is possible to use different $\mathbf{k}/\mathbf{q}$ coarse and fine grids for the two spin channels since there is no guarantee that both spin channels would behave equivalently. Nevertheless, in our case, we find that convergence is achieved with the same settings. A good compromise between accuracy and computational cost is obtained with a $8^{3}$ $\mathbf{k/q}$ coarse grid; see Supplementary Material S7 [83] for further discussion.
The computed resistivity of Fe is presented in Fig. 6(a) and shows that the BTE with electron-phonon scattering explains only partially the resistivity and tends to underestimate it. We find that PBE results are slightly closer to experiment due to the improved phonon dispersion, see Fig. 2(c), and a small decrease in Fermi velocity; see Supplementary material S8 [83]. We also find that the SERTA is a good approximation to the iterative BTE solutions, which is typical for non-magnetic metals [94]. Our results are qualitatively similar to a previous study using the Ziman formula [49] where quantitative agreement with the experimental results is obtained at low temperatures. However, we notice that the electron-phonon contribution to resistivity does seem to dominate at low temperature, see inset in Fig 6(a), and at temperatures up to 300 K our calculations suggest that electron-phonon scattering can account for at least 75% of the experimental resistivity when using the PBE functional. For higher temperatures, the observed underestimation indicates that electron-phonon scattering is not the sole mechanism contributing to the resistivity.
@@
We continue our analysis by investigating the case of ferromagnetic Ni. As seen in Fig. 6(b), the computed resistivity for Ni is also underestimated in the entire temperature range compared to experiment. However, above 350 K the resistivity originating from electron-phonon scattering is almost one order of magnitude less than the experimental resistivity. In addition, the agreement at low temperatures is worst than in Fe since there is an important underestimation of the resistivity already in the 50-100 K temperature range, see the inset in Fig. 6(b). Interestingly, we note that using the PBE functional does not improve significantly in spite of the remarkably better agreement of the phonon dispersion with experiments, see Fig.2(d). This hints at an important distinction with respect to Fe.
@@
In ferromagnetic compounds, contributions to the resistivity originating from magnons or spin fluctuations are expected to be significant [98, 99]. One of the reasons why magnons may explain the missing resistivity of Ni and Fe above 300 K is due to their energy scale which is around 8 to 9 times bigger than the phonons [100]. The magnon population scales as $(T/T_{\rm Curie})^{2}$ [99], and will be smaller in Fe than in Ni for a given temperature due to the lower Curie temperature ($T_{\rm Curie}$) of 627 K  [96], compared to 1043 K in Fe [97]. Consequently, the saturation that gives a linear relationship with the phonons occurs at significantly elevated temperatures, and the “low temperature” $T^{2}$ behavior of the magnons is observed over a larger range of temperatures. Although this explains the increase in the importance of magnon at high temperature, it does not explain the marked difference between Fe and Ni.
@@
In fact, the magnon population is crucially dependent on the density of states at the fermi level of the opposite spin channel [101]. The large density of states of the spin down (minority) channel in Ni (see Fig. 4(b)) above the Fermi level allows for a large magnon creation by converting spin up (majority) electrons into spin down (minority) electrons. In the case of Fe, magnon creation is suppressed by the small accessible phase space of minority electrons [101]. In addition to the difference in magnon creation, our calculations show that the spin up electrons in Ni experience far less scattering from electron-phonon, and, as a result, our calculations show that $\rho_{\alpha\beta}\approx\rho^{\uparrow}_{\alpha\beta}$ with an order of magnitude difference in the conductivities $\sigma_{\alpha\beta}^{\uparrow}\gg\sigma^{\downarrow}_{\alpha\beta}$, while for Fe $\rho^{\uparrow}_{\alpha\beta}\sim\rho_{\alpha\beta}^{\downarrow}$. Since the spin up channel for Ni shows a larger lifetime broadening due to magnons [101], the total lifetime $\tau_{n\mathbf{k}}^{\uparrow}$ is smaller than with electron-phonon scattering only, explaining the underestimated resistivity compared to experiment in Ni.
In summary, comparing the calculated resistivity at low temperature with the experimental data of Fe and Ni, we find that the behavior of these two compounds is fundamentally different: magnons dominate the scattering in Ni while electron-phonon is the dominant scattering mechanism in Fe up to room temperature.
@@ V.2 Role of magnetism on physical properties
One of the key aspects of including magnetism is the fact that the electronic structure of the compounds changes with magnetism, as does the phonon dispersion. As seen in Fig. 7(a), Fe is not dynamically stable when we use the non-magnetic (NM) approximation. This is due to a fermionic pressure that increases the lattice parameters of Fe when magnetism is allowed [102, 103, 104]. The instabilities found in the phonon dispersion for a BCC structure in the NM approximation are known to be related to distortions that lead to hexagonal phases [105]. This is also consistent with the fact that in the crystallographic phase diagram of Fe, we find a hexagonal phase at pressures a bit above 10 GPa [106, 107]. It is also reported that the FM phase becomes paramagnetic, antiferromagnetic or even diamagnetic at sufficiently high pressures.
Interestingly, the phonon dispersion for Ni does not change much by using the NM approximation; see Fig. 7(b). In fact, one can notice that the FM phase of Ni produces only a small overall imbalance in the spin occupations since the experimental magnetic moment per unit cell is $M_{\mathrm{exp}}=0.616~\mu_{\mathrm{B}}$ [108] when compared with Fe that presents a magnetic moment of $M_{\mathrm{exp}}=(2.2\pm 0.1)~\mu_{\mathrm{B}}$ [109, 110, 111]. Thus, one may wonder how different the resistivity in the NM phase is, when compared with that in the FM collinear case for Ni. In that regard, we repeat the calculations with the BTE for the resistivity of Ni with PBE (see LDA in the supplementary material S10 [83]) but using the NM phase and we compare it in Fig. 8(b) with the FM case and with the experimental data.
@@
Although the limitation of the NM approximation has been highlighted previously [112, 113, 114, 115, 116], it is common to find studies that still use the NM approximation to model spin-disordered paramagnetic, the ferromagnetic or antiferromagnetic phases. For this reason, we also compute the resistivity of NM Fe as a function of temperature and present it in Fig. 8(a). As expected, we obtain an extremely overestimated resistivity when compared to the FM one since we neglect scattering originating from soft modes. Importantly, we show that even though the phonons are similar in Ni, neglecting the spin degree of freedom yields the wrong conclusion that the resistivity is dominated by electron-phonon interactions. We remark that, in spite of the seemingly better agreement with the experiment, the NM results actually overestimate the experimental data, which should never happen since we are not considering the full set of scattering channels. Moreover, the temperature-dependent trend of the NM result is inconsistent with experimental observation.

### v2

@@ III Methods
DFT calculations are performed using Quantum ESPRESSO (QE[41, 40] version 7.4 and Abinit version 10.2.5 [44, 45, 46, 112]. Scalar-relativistic norm-conserving pseudopotentials [53] from the standard accuracy table of the PseudoDojo project [108] are used. We use the local density approximation (LDA) [86] and the generalized gradient approximation (GGA) with the Perdew Burke Ernzerhof (PBE) functional [87]. The interpolated phonon dispersions and dynamical matrices are obtained with DFPT [9, 98, 47, 8]. The maximally localized Wannier functions are obtained using the wannier90 package version 3.1 [88]. Direct evaluation of spin-resolved EPME is obtained by modifying QE to print the values. We provide the modified code on the Materials Cloud Archive [20] along with all data to reproduce the results presented here. The interpolation of electronic band structures, phonon band structures, and electron phonon matrix elements, as well as the calculation of transport properties and electron-phonon coupling matrix elements are performed by modifying the EPW package version 5.9 [69]. This feature will be released in a future EPW version and we also provide the code on the Materials Cloud Archive [20].
The cutoff energy for the plane wave expansion, $\bf k$-meshes for the electrons, and the $\bf q$-meshes for the phonons are selected so that an arbitrary increase in those parameters produces a total energy change of less than 1 meV per unit cell. This corresponds to 100 Ry for both Fe and Ni, as well as a $\bf k$-point grid of 243 and a $\bf q$-point grid of 83 for both Ni and Fe. Structural relaxations are performed with a convergence criterion of $\Delta E<10^{-12}$ Ry, and with a pressure of less than 1 bar. In the case of phonons, the convergence criteria in order to obtain accurate force constants was tr2_ph $10^{-20}$ for QE while for Abinit the convergence criteria is on the energy $\Delta E<10^{-20}$ Ha. We converge the resistivity below 1% above 150 K using EPW, and validate the results with an in-house developed version of the ElectronPhonon.jl package [70, 72], which relies on the real-space EPME computed with the EPW code [91, 69] that we extend to the collinear magnetic case for the purpose of this work. Excellent agreement is obtained between the two codes, see Supplementary Material S2 [20] for more information.
@@ IV Electron-phonon coupling with magnetism
In this work, we study BCC Fe and FCC Ni, both in their ferromagnetic configuration using the collinear approximation. We also consider non-magnetic Pb and verify that the LSDA implementation yields the same results as the LDA case as described in Supplementary Material S3 [20].
@@ IV.1 Verification of the direct evaluation of the electron-phonon coupling
To verify our implementation, we compare the QE results with those obtained with Abinit. Although both codes match well in non-magnetic compounds [90, 16, 93], no validation study had yet been performed in the magnetic case. To this end, we compare basic quantities including total energy, total magnetization and phonon frequencies while keeping the computational parameters as close as possible. We find an excellent level of agreement between the two codes that is similar to the non-magnetic case, see Supplementary Material S4 [20].
@@
We compare the absolute value of the deformation potentials between the two codes. To achieve this, we note that QE performs an average over all degenerate states in electronic and phonon band indices. We therefore applied the same average to the Abinit results and present a comparison for ferromagnetic Fe in Fig. 1 between the two codes for $\mathbf{k}=\bm{\Gamma}$ and bands $m=n=5$. We find a good agreement between the two codes with a maximum error of less than 0.02 meV/bohr. We provide the same comparison for Ni in Supplemental Material S5 [20] where we find an error below 0.01 meV/bohr.
@@ IV.2 Verification of the electron phonon interpolation
Figure 3: Spin resolved deformation potential for the three phonon branches $\nu$ at $\mathbf{k}=\Gamma$ and as a function of wave vector $\mathbf{q}$. The color scheme is presented as an inset in the figure $\uparrow$ (red/orange/yellow/green) and $\downarrow$ (blue/pink/purple). The solid dots are direct evaluation of the deformation potential with Quantum ESPRESSO. The lines are the EPW interpolated deformation potential from $4^{3}$, $6^{3}$, $8^{3}$, $10^{3}$ and $12^{3}$ $\mathbf{k/q}$ homogeneous coarse grids with the colors provided in the inset of the figure. The electronic bands $m$ and $n$ are summed using Eq. (6) from second to sixth lowest electronic bands shown in Fig. 2(a) and 2(b) for (a) Fe and (b) Ni.
We first compare in Fig. 2(a) and 2(b) the direct and Wannier-interpolated electronic band structures of ferromagnetic Fe and Ni respectively, obtaining an energy mismatch of less than 0.1 meV for all the displayed $\mathbf{k}$-points and bands. This shows that both spin channels are correctly reproduced by the interpolation procedure. We then compute in Fig. 2(c) and 2(d) the phonon dispersion of ferromagnetic Fe and Ni respectively using LDA and PBE. Overall, LDA overestimates by about 10% the phonon frequency with respect to neutron scattering experiments [14, 18]. Using the PBE functional yields a much better agreement, in line with a previous study [29]. Moreover, we validate the interpolated electron-phonon deformation potential by comparing to direct calculations in Fig. 3, for electronic bands crossing the Fermi level and at the electronic zone center. We show the two spin channels for Fe and Ni, and find that convergence is achieved for $\times12$$\times$12 $\mathbf{k}/\mathbf{q}$ coarse grids in both cases. For such grids, we find excellent agreement between interpolated and independently calculated reference direct calculations. In addition, we also compare the EPW interpolation with the Abinit interpolation finding similar results; see Supplementary Material S5 [20]. Finally, we verify that the Hamiltonian, dynamical matrices, and EPME all decay rapidly in real space; see Supplementary Material S6 [20].
@@ V Physical properties and electron-phonon quantities
Compound Magnetism $\lambda$ $\lambda^{\uparrow}$ $\lambda^{\downarrow}$
Fe (BCC) FM 0.2961 0.2180 0.0781
Ni (FCC) FM 0.2767 0.0040 0.2727
Table 1: Total and spin-resolved electron phonon coupling strength $\lambda$ for ferromagnetic (FM) Fe and Ni.
@@
We also compute the Eliashberg spectral functions and show the result in Fig. 5. We see in Fig. 5(a) that the Fe spin majority channel is the dominant contribution for most of the energy range. In contrast, for Ni the spin minority dominates throughout and the spin majority is negligible, see Fig.  5(b). This important difference is due to the density of states at the Fermi level in Ni, see Fig. 4. Recently, some authors have explored phonon-mediated superconductivity in magnetic systems by means of the mass enhancement $\lambda$ and the Eliashberg spectral function [121]. Although the validity of using the Allen-Dyne formula [3] in magnetic materials has not yet been demonstrated, we report superconducting results for the non-magnetic and magnetic cases in Supplementary Material S7 [20]. In both cases and for both materials, we find a vanishing superconducting temperature in agreement with experimental observation.
@@ V.1 Resistivity of magnetic metals
We now focus on the resistivity of metals and solve the BTE for both spin channels in ferromagnetic Fe and Ni. We note that when converging the resistivity and in general electron-phonon related quantities, it is possible to use different $\mathbf{k}/\mathbf{q}$ coarse and fine grids for the two spin channels since there is no guarantee that both spin channels would behave equivalently. Nevertheless, in our case, we find that convergence is achieved with the same settings. A good compromise between accuracy and computational cost is obtained with a $8^{3}$ $\mathbf{k/q}$ coarse grid; see Supplementary Material S8 [20] for further discussion.
The computed resistivity of Fe is presented in Fig. 6(a) and shows that the BTE with electron-phonon scattering explains only partially the resistivity and tends to underestimate it. We find that PBE results are slightly closer to experiment due to the improved phonon dispersion, see Fig. 2(c), and a small decrease in Fermi velocity; see Supplementary material S9 [20]. We also find that the SERTA is a good approximation to the iterative BTE solutions, which is typical for non-magnetic metals [71]. Our results are qualitatively similar to a previous study using the Ziman formula [73] where quantitative agreement with the experimental results is obtained at low temperatures. However, we notice that the electron-phonon contribution to resistivity does seem to dominate at low temperature, see inset in Fig 6(a), and at temperatures up to 300 K our calculations suggest that electron-phonon scattering can account for at least 75% of the experimental resistivity when using the PBE functional. For higher temperatures, the observed underestimation indicates that electron-phonon scattering is not the sole mechanism contributing to the resistivity.
@@
We continue our analysis by investigating the case of ferromagnetic Ni. As seen in Fig. 6(b), the computed resistivity for Ni is also underestimated in the entire temperature range compared to experiment. However, above 350 K the resistivity originating from electron-phonon scattering is almost one order of magnitude less than the experimental resistivity. In addition, the agreement at low temperatures is worse than in Fe since there is an important underestimation of the resistivity already in the 50-100 K temperature range, see the inset in Fig. 6(b). Interestingly, we note that using the PBE functional does not improve significantly in spite of the remarkably better agreement of the phonon dispersion with experiments, see Fig.2(d). This hints at an important distinction with respect to Fe.
@@
In ferromagnetic compounds, contributions to the resistivity originating from magnons or spin fluctuations are expected to be significant [96, 116]. One of the reasons why magnons may explain the missing resistivity of Ni and Fe above 300 K is due to their energy scale which is around 8 to 9 times bigger than the phonons [52]. The magnon population scales as $(T/T_{\rm Curie})^{2}$ [116], and will be smaller in Fe than in Ni for a given temperature due to the lower Curie temperature ($T_{\rm Curie}$) of 627 K  [67] in Ni, compared to 1043 K in Fe [38]. Consequently, while at 300 K the phonon contribution to the resistivity would sow a linear “high temperature behavior” , the magnon contribution would still be within its “low temperature regime” with a $T^{2}$ behavior. Although this explains the increase in the importance of magnons at temperatures beyond 200-300 K, it does not explain the marked difference between Fe and Ni.
@@
In fact, the magnon population is crucially dependent on the density of states at the Fermi level of the opposite spin channel [83]. The large density of states of the spin down (minority) channel in Ni (see Fig. 4(b)) above the Fermi level allows for a large magnon creation by converting spin up (majority) electrons into spin down (minority) electrons. In the case of Fe, magnon creation is suppressed by the small accessible phase space of minority electrons [83]. In addition to the difference in magnon creation, our calculations show that the spin up electrons in Ni experience far less scattering from electron-phonon, and, as a result, our calculations show that $\rho_{\alpha\beta}\approx\rho^{\uparrow}_{\alpha\beta}$ with an order of magnitude difference in the conductivities $\sigma_{\alpha\beta}^{\uparrow}\gg\sigma^{\downarrow}_{\alpha\beta}$, while for Fe $\rho^{\uparrow}_{\alpha\beta}\sim\rho_{\alpha\beta}^{\downarrow}$. Since the spin up channel for Ni shows a larger lifetime broadening due to magnons [83], the total lifetime $\tau_{n\mathbf{k}}^{\uparrow}$ is smaller than with electron-phonon scattering only, explaining the underestimated resistivity compared to experiment in Ni. In summary, comparing the calculated resistivity at low temperature with the experimental data of Fe and Ni, we find that the behavior of these two compounds is fundamentally different: magnons dominate the scattering in Ni while electron-phonon is the dominant scattering mechanism in Fe up to room temperature. We also consider the case of ferromagnetic Co HCP finding that the resistivity would be a case in between Fe and Ni; see Supplementary material S10 [20]
@@ V.2 Role of magnetism on physical properties
One of the key aspects of including magnetism are the changes on both electronic and phonon band dispersion. Therefore, we consider the effects of using the non-magnetic (NM) approximation in Fe and Ni by relaxing their BCC and FCC structures and calculating their respective phonon dispersions, see Fig. 7. As we can see in Fig. 7(a), Fe is not dynamically stable when we use the NM approximation. This is due to a fermionic pressure that increases the lattice parameters of Fe when magnetism is allowed [60, 80, 75, 64]. These instabilities found in the phonon dispersion for a BCC structure in the NM approximation are known to be related to distortions that lead to hexagonal phases [49]. This is also consistent with the fact that in the crystallographic phase diagram of Fe, we find a hexagonal phase at pressures a bit above 10 GPa [28, 103].
Interestingly, the phonon dispersion for Ni does not change much by using the NM approximation; see Fig. 7(b). In fact, one can notice that the FM phase of Ni produces only a small overall imbalance in the spin occupations since the experimental magnetic moment per unit cell is $M_{\mathrm{exp}}=0.616~\mu_{\mathrm{B}}$ [30] when compared with Fe that presents a magnetic moment of $M_{\mathrm{exp}}=(2.2\pm 0.1)~\mu_{\mathrm{B}}$ [107, 85, 7]. Thus, one may wonder how different the resistivity in the NM phase is, when compared with that in the FM collinear case for Ni. In that regard, we repeat the calculations with the BTE for the resistivity of Ni with PBE (see LDA in the supplementary material S11 [20]) but using the NM phase and we compare it in Fig. 8(b) with the FM case and with the experimental data.
@@
Although the limitation of the NM approximation has been highlighted previously [110, 109, 118, 119, 74], it is common to find studies that still use the NM approximation to model spin-disordered paramagnetic, the ferromagnetic or antiferromagnetic phases [15, 78, 50]. In addition, some studies have also neglected the soft modes and still perform transport calculations [120, 48, 100] For this reason, and in order to highlight how bad would be to perform this calculations, we also compute the resistivity of NM Fe as a function of temperature and compare it with the FM case in Fig. 8(a). The NM calculation with negative phonon frequencies yield an extremely overestimated resistivity when compared to the FM one and the experimental results. Importantly, we show that even though the phonons are similar in Ni, neglecting the spin degree of freedom yields the wrong conclusion that the resistivity is dominated by electron-phonon interactions. We remark that, in spite of the seemingly better agreement with the experiment, the NM results actually overestimate the experimental data, which should never happen since we are not considering the full set of scattering channels. Moreover, the temperature-dependent trend of the NM result is inconsistent with experimental observation.